OpenStudy (anonymous):

PLEASE PLEASE HELP ME Describe the differences between the graph of y = -3(x + 7)2 - 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

Did you mean y = -3(x + 7)^2-10 and y=x^2 ?

OpenStudy (anonymous):

Differences: a) the vertex of eq.1 is at (-7,-10) while the vertex of eq.2 is at (0,0) b) eq.1 opens downward and eq.2 opens upward

OpenStudy (anonymous):

@edward_trex Yes! sorry I didn't catch that! thank you so much for helping me